Child Health and Development Studies (CHDS) has been collecting data about expectant mothers in Oakland, CA since 1959. One of the measurements taken by CHDS is the weight increase (in pounds) for expectant mothers in the second trimester. In a fictitious study, suppose that CHDS finds the average weight increase in the second trimester is 14 pounds. Suppose also that, in 2015, a random sample of 43 expectant mothers have mean weight increase of 16.2 pounds in the second trimester, with a standard deviation of 5.7 pounds. A hypothesis test is done to see if there is evidence that weight increase in the second trimester is greater than 14 pounds. Find the p -value for the hypothesis test. The p -value should be rounded to 4 decimal places.
Answer #1

Solution :

This is the one tailed test .

The null and alternative hypothesis is ,

H0 :  \mu = 14

Ha : \mu > 14

Test statistic = z

= (\bar x - \mu) / \sigma / \sqrtn

= (16.2 - 14) / 5.7 / \sqrt43

Test statistic = 2.53

P(z > 2.53) = 1 - P(z < 2.53) = 1 - 0.9943

P-value = 0.0057‬